# Exploring and Benchmarking the Planning Capabilities of Large Language Models

### v1

@@ Abstract
We seek to elevate the planning capabilities of Large Language Models (LLMs)investigating four main directions. First, we construct a comprehensive benchmark suite encompassing both classical planning domains and natural language scenarios. This suite includes algorithms to generate instances with varying levels of difficulty, allowing for rigorous and systematic evaluation of LLM performance. Second, we investigate the use of in-context learning (ICL) to enhance LLM planning, exploring the direct relationship between increased context length and improved planning performance. Third, we demonstrate the positive impact of fine-tuning LLMs on optimal planning paths, as well as the effectiveness of incorporating model-driven search procedures. Finally, we investigate the performance of the proposed methods in out-of-distribution scenarios, assessing the ability to generalize to novel and unseen planning challenges.
Exploring and Benchmarking the
Planning Capabilities of Large Language Models
Bernd Bohnetthanks: Equal contribution                        Azade Nova11footnotemark: 1                        Aaron T Parisi                        Kevin Swersky                        Katayoon Goshvadi
Hanjun Dai                        Dale Schuurmans                        Noah Fiedel                        Hanie Sedghi
Google DeepMind
@@ 1 Introduction
Intelligent agents require the ability to plan— to proactively chart a course of action to achieve their objectives. This capacity for strategic foresight is fundamental to intelligent behavior (Russell and Norvig, 2016).
While classical search algorithms have long been the cornerstone of planning studies, machine learning techniques, particularly Monte-Carlo Tree Search (MCTS) and reinforcement learning, have emerged as useful additions, significantly expanding the capabilities of modern planning systems.
#  Planning Domain Definition Language
(define (problem BW-rand-3)
(:domain blocksworld-4ops)
(:objects A B C)
  (:init
@@
# Planning: LLM generates a PDDL plan.
@@
The goal is: C is on B. A is on C.
# Planning: LLM generates natural language plan
@@
Figure 1: Blocksworld planning in PDDL and natural language.
With the advent of powerful large language models (LLMs), there are new opportunities to both revisit classical planning problems, and to further explore new problems through natural language specification that reflects the ambiguity and uncertainty of real-world domains. Planning capability is important for many tasks such as meeting scheduling and trip planning. Moreover, it is a crucial skill for assisting users with task planning, text planning, or performing tasks in the real world. Research is already underway to leverage the commonsense knowledge of LLMs in real-world tasks (Huang et al., 2022; Singh et al., 2023; Ding et al., 2023) and to generate plans (Valmeekam et al., 2023; Hao et al., 2023; Guan et al., 2024). This research has shed some light on LLMs’ struggle with planning tasks. They may produce ineffective or even incorrect plans, even in straightforward scenarios.
This paper focuses on improving the planning capability of LLM systems through focusing on both instructing the model and enhancing the search strategy required for planning. We provide a novel exploration of in-context learning(ICL), utilizing the emerging paradigm of many-shot learning with long context. We compare this approach with traditional fine-tuning methods and investigate both in-domain and out-of-domain generalization.
Specifically, we explore two distinct planning representations: the formal Planning Domain Definition Language (PDDL) McDermott et al. (1998), which provides a standardized representation for classical planning problems and allows for rigorous plan validation; and natural language, which offers a more flexible and intuitive representation better reflecting real-world scenarios. For both scenarios, we provide a code for generating as many instances with a degree of difficulty of choice. We also provide a mapping method for translating PDDL benchmarks to natural language and measure the performance of the generated benchmarks. The generated planning tasks are scalable and can grow to examine and assist stronger models.
@@
We observe that (1) carefully instructing the model using ICL leads to a significant boost in planning performance, which can be further improved by using a many-shot approach in long context. (2) finetuning with the optimal plan can lead to near-perfect accuracy for models that are much smaller than SOTA LLMs. (3) The proposed plans show generalization to unseen instances of the environment: instances with similar complexity have the same accuracy, (4) For ICL, teaching easier instances leads to better performance on hard instances compared to teaching hard instances. (4) Finetuning leads to better performance for both in-domain and out-of-domain scenarios, even with a much smaller model, compared to ICL with long-context on the SOTA LLM models. (5) Integrating the model with a search procedure such as MCTS can help earlier versions and smaller models perform closer to SoTA foundational models.
Reasoning and planning are critical cognitive skills in humans. These abilities can be articulated through natural language; reasoning has been explored in natural language processing (NLP) Bauer and Bansal (2018); Ning et al. (2020); Yu et al. (2020), whereas planning has not yet been very much represented within the NLP research community except in the context of text planning Reiter and Dale (2000); Gatt and Krahmer (2018).
In summary, our primary contributions are:
[1] We provide a scalable benchmark suite in both PDDL and natural language to measure planning capability of LLMs.
[2] We teach planning capability to LLMs using both in-context learning (and recent many-shot paradigm) and fine-tuning strategies as well as investigate the effect of better search mechanisms.
[3] We investigate plan generalization capability of our methods.
The rest of the paper is organized as follows. We describe the benchmark suite and approaches to enhance LLMs planning capability in Section 2; Describe experiments setting and discuss the results in Section 3 and conclude in Section 4. The limitations and ethics statements are in Section 56 respectively.
@@ 1.1 Related Work
There has been some investigations of planning capabilities of LLMs stating that these models struggle with solving planning tasks (Hao et al., 2023; Valmeekam et al., 2023, 2024), In contrast, we show that LLMs, once instructed well, are capable of solving such tasks and one can reach close to accuracy of one for some scenarios. We run experiments on the same BlocksWorld benchmark, and on more difficult cases (more number of blocks) and similar to these works we use verifiers to compute accuracy. See Section 3.4 for details.
Xie et al. (2024) proposed a TravelPlanner benchmark and showed GPT4-Turbo can solve some of the benchmark tasks with a success rate of $0.6\%$. The TripPlanner benchmark used in our work has two main differences, one is that it is not an agent environment and rather a natural language benchmark, the other is that it has unique answers due to carefully designed constraints. Therefore, we chose the latter for our experiments.
Stechly et al. (2024) suggested that LLMs are not capable of generalizing their plans on BlocksWorld if one uses chain-of-thought (CoT) (Wei et al., 2022). In this work we show positive results in terms of generalization performance.
@@
In addition to standard few-shot prompting, we also make use of search procedures tree-of-thought (ToT) (Yao et al., 2023) and Monte-Carlo Tree Search (MCTS) (Hao et al., 2023). Each of these expand on conventional chain-of-thought reasoning. We demonstrate that these methods can considerably improve an LLMs planning capabilities, beyond ICL and intrinsic chain-of-thought.
2 Planning Datasets and Methods
To evaluate the planning capabilities of large language models (LLMs), we constructed a benchmark suite for planning. On these benchmarks, we assessed LLM performance using In-Context Learning (ICL), Supervised Fine-Tuning (SFT), and Search-based Planning.
PDDL (Planning Domain Definition Language): PDDL McDermott et al. (1998) is a standardized language used in artificial intelligence for representing planning problems. PDDL provides a way to describe the initial state of the world, the desired goals, the actions that can be taken, and the effects of those actions. PDDL has two main components (1) Domain: Describes the general characteristics of the planning problem, including the types of objects, actions, and predicates (conditions that can be true or false). (2) Problem: Defines a specific instance of the planning problem within the domain, including the initial state of the world and the goals to be achieved.
We selected three datasets that use PDDL, constructed subsets of them, and additionally mapped these datasets to natural language for an additional evaluation task. Additionally, we selected two native natural language datasets, containing Trips-Planning and Calendar-Scheduling tasks Zheng et al. (2024). For the PDDL-based datasets, we selected BlocksWorld, Logistics, and Mini-Grid. We translated the PDDL problem descriptions of all these datasets into natural language to compare the performance of problems provided as formal and informal problem description.
In Section 2.1, we describe the creation of the PDDL datasets and the generation of additional natural language tasks from these PDDL problem definitions. Note that our generation code can be used to generate any number of instances from these benchmarks with arbitrary complexity (through the addition of additional constraints or elements). For all these PDDL-based tasks, we calculate accuracy of the plan generated by LLM using a verifier tool (Fox and Long, ). In Section 2.2, we describe the native Natural Language Benchmarks.
2.1 Parallel PDDL and Natural Language Benchmarks
@@
This procedure is iterated with increasingly difficult configurations for a selected number of problems. The result of this procedure are datasets that comprise a set of problems in PDDL and their solutions, i.e., the plans to solve the problems.
Importantly, this procedure enables us to create datasets with increasingly difficult problems, which are appropriate for assessing the ability to plan using different methods, such as in-context learning versus Supervised Fine-Tuning. Moreover, we can scale the dataset generation and create as many instances as needed for different investigations. We provide the details of our benchmark suite below.
@@
BlocksWorld
BlocksWorld is a standard planning problem from International Planning Conference (IPC)-2000  [^2] https://github.com/potassco/pddl-instances/tree/master/ipc-2000. This domain consists of a set of blocks, a table and a robot hand, where the goal is to move from one block configuration to another.
Algorithm 1 Create BlocksWorld Dataset
function CreateDatasetBW($num\_blocks$, $n$)
     $dataset\leftarrow[]$triangleright Initialize an empty list
     for $problem\_id\leftarrow 1$ to $n$ do
         $b\leftarrow$ RandomUniform($3$, $num\_blocks$)
         $initStacks\leftarrow$ CreateStacks($b$)
         $goalStacks\leftarrow$ CreateStacks($b$)
         if $initStacks==goalStacks$ then
              continue $\triangleright$ Skip equal stacks.
         end if
         $problem\leftarrow$CreatePro($initStacks,goalStacks$)
         $plan\leftarrow$FastDownward($problem,domain$)
         $dataset\leftarrow dataset+[(problem,plan)]$
     end for
     return $dataset$
end function
We generate a datase for 3 to 7 blocks. We produced 28k unique BlocksWorld samples. From these, 25.5k were randomly selected for the training set and 2,500 for the validation set. We describe the mapping via templates from PDDL to natural language in the appendix B.2.
Logistics
Logistics is an AI planning problem from IPC-1998  [^4] https://github.com/potassco/pddl-instances/tree/master/ipc-1998 expressed in PDDL that involves arranging the delivery of packages to their destinations using trucks within cities and airplanes between cities. The aim is to optimize transportation modes under constraints such as vehicle capacities and locations, showcasing model’s ability to manage multi-step logistics efficiently.
Mini-Grid
Mini-Grid is a task from Artificial Intelligence Planning Systems (AIPS)-1998  [^5] https://github.com/AI-Planning/pddl-generators/tree/main/minigrid, also expressed in PDDL. We create various floorplans with rooms containing random configurations of key shapes. The goal then is for a robot to navigate from an initial position to a designated goal cell.
@@ 2.2 Native Natural Language Planning Benchmarks
In addition to the PDDL-derived planning tasks, we also rely on several natural-language benchmarks.
Trip Planning
Trip Planning is a task from NaturalPlan (Zheng et al., 2024) benchmark focusing on planning a trip itinerary under given constraints. The goal of the task is to find an itinerary satisfying constraints such as the order of visiting N cities. It includes enough constraints for each instance such that there is only one solution to the task, which makes the evaluation of the predictions straightforward.
Calendar Scheduling
Calendar Scheduling from the NaturalPlan (Zheng et al., 2024) benchmark represents the task of scheduling a meeting of either 30 minutes or an hour among up to 7 attendees. The attendees may have a busy schedule or a light schedule with less than half of the working hours spent in meetings.
2.3 Methods
In-Context Learning
For in-context learning (ICL) (Brown et al., 2020), we adhere to the standard procedure, employing a prompt containing several examples for the task. Each example comprises a planning problem statement and its corresponding solution, referred to as a shot. Following the examples, the test problem is added without the solution which completes the prompt. Subsequently, a LLM receives this prompt and is expected to generate a plan following the format and logic of the examples in the prompt. See Appendix A for examples of prompts.
In this work, we explore potential gains by moving beyond few-shot learning and utilizing a large number of shots (Agarwal et al., 2024).
Planning with Search Procedures
In addition to standard few-shot in-context methods, we explore two natural-language search strategies: Monte Carlo Tree Search (MCTS)(Hao et al., 2023), and Tree-of-Thought (ToT). (Yao et al., 2023) proposed Tree-of-Thought, expanding the notion of chain-of-thought (CoT) (Wei et al., 2022) reasoning by considering a chain-of-thought a single path through a tree constructed by branching at each step of the search procedure. For reasoning tasks each node represents the current state of the reasoning chain, which can be expanded by branching children off of the node in question.
Building on tree-of-thought, (Hao et al., 2023) modified the Monte Carlo Tree Search (MCTS) procedure described in previous works (Silver et al. (2017)) to work with LLMs in place of the policy and world models. Similar to ToT, the MCTS procedure at each node represents a specific state of the reasoning chain. However, it has two key differences from the ToT procedure: 1) it uses a function of the Q-value to drive exploration and exploitation based on node selection, and 2) beyond the ’state’ of the reasoning chain, an additional ’world state’ is maintained, allowing tasks which interact with an environment to approximate and reason about the environment explicitly in the reasoning tree.
Supervised Fine-Tuning
Supervised fine tuning (SFT) (Ouyang et al., 2022) has proved to be an effective method for teaching LLMs various capabilities. We investigate impact of SFT on teaching planning to LLMs. We use the Fast-Downward classical planner to generate the optimal plan and investigate the effect of fine tuning on that on the planning performance for two different benchmarks with different levels of difficulty.
Plan generalization
As for all machine learning capabilities, it is desired for models to generalize to unseen inputs. Therefore, we investigate plan generalization capabilities of LLMs. Plan generalization has three main categories: (1) Generalize to unseen instances of the same environment (2) Generalize to renaming of actions and objects (3) Generalize to unseen plan environments. Here, we focus on the first category which sits at the core of the desired capability. This is equivalent to Out of distribution (OOD) scenarios in machine learning. For this investigation we look into effect of both SFT as well as ICL on the generalization performance for two benchmarks.
@@ 3 Experiments
As mentioned earlier, without intervention LLMs often struggle with even simple planning tasks. They often lack the information on how to structure their plan constructively, and with respect to constraints in the task. In this section we present our experimental results for improving LLMs planning capability; and investigate plan generalization (i.e., the ability to generalize to unseen instances) in several scenarios.
For PDDL experiments, we measure accuracy of the generated plan with a verifier. For natural language experiments we rely on either recasting the task to PDDL and verifying with a verifier (Fox and Long, ), or extracting the answer and comparing it to expected results (Zheng et al., 2024).
In all experiments, GPT-4 refers to GPT-4 Turbo and we omit "Turbo" for space constraints.
3.1 In-context learning
@@
Figure 2: PDDL Planning and Natural Language Planning with few-shots. Natural language text are generated from formal PDDL problem definitions.
In this section, we report performance of models on both classical planning benchmarks expressed in the standard Planning Domain Definition Language (PDDL) as well as in natural language. See Appendix A for examples of prompts for all planning tasks considered in this section.
We evaluate the planning capability of the model as we add more examples (“shots") into the context, inspired by the success of many-shot learning across a large number of tasks (Agarwal et al., 2024). The challenge of “in-context planning" involves understanding a specific task and problem through a limited number of examples. Additionally, it requires the models to produce a solution without checking each planning step to confirm if a proposed move is correct. The model has to create a plan in a single inference step, keeping ’in mind’ all the constraints the task imposes.
Figure 2 shows the in-context learning performance on classical planning and natural language benchmarks as we vary the number of shots. Overall, we notice that for natural language and PDDL scenarios, models have similar trends in terms of planning accuracy as we increase the number of shots. Moreover, different models are impacted differently as we provide additional number of shots, e.g., Gemini 1.5 Pro outperforms other models indicating that the model not only can plan better with a fewer number of examples/shots, it can also make effective use of additional and longer context. Gemini 1.5 Flash -a smaller, faster and more efficient model than Gemini 1.5 Pro is still outperformed by Gemini 1.5 Pro but occasionally matching GPT-4 Turbo performance. The details of results for each tasks are provided below.
BlocksWorld
We generated BlocksWorld problem instances of 3 to 7 blocks for in-context learning with the method described in Appendix B.1.
Figure 2(a)2(d) show the performance of Gemini 1.5 models on this benchmark as we increase the number of few-shot examples. We note that as we increase the number of shots GPT-4 Turbo’s performance increases while Gemini 1.5 Pro’s performance saturates or degrades as we go beyond 40 shots. The 1-shot planning capability of Gemini 1.5 Pro and Gemini 1.5 Flash reaches reaches 35% and 26%, while GPT-4 Turbo performance is close to zero. Moreover the 40-shots planning capability of Gemini 1.5 Pro reaches 48% range which performs better than the best (200-shots) performance of GPT-4 Turbo, which peaks at 43%.

### v2

# Exploring and Benchmarking Planning Capabilities of Large Language Models

@@ Abstract
Classical and natural language planning tasks remain a difficult domain for modern large language models (LLMs). In this work, we lay the foundations for improving planning capabilities of LLMs. First, we construct a comprehensive benchmark suite encompassing both classical planning benchmarks and natural language scenarios. This suite includes algorithms to methodically generate instances of tasks with varying levels of difficulty, allowing for rigorous and systematic evaluation of LLM performance. Next, we investigate the use of many-shot in-context learning to enhance LLM planning, exploring the relationship between increased context length and improved planning performance. In addition, we demonstrate the positive impact of fine-tuning LLMs on optimal planning paths. We also probe the efficacy of chain-of-thought reasoning methods to improve LLM planning performance. Moreover, we probe the performance of the proposed methods in out-of-distribution scenarios, assessing the ability to generalize to novel and unseen planning challenges. Finally, we investigate model’s failure modes and reveal insights that hold true across different benchmarks.
@@ 1 Introduction
Intelligent agents require the ability to plan to proactively chart a course of action to achieve their objectives. This capacity for strategic foresight is considered fundamental to intelligent behavior (Russell & Norvig, 2016). While classical search algorithms have long been the cornerstone of planning studies, machine learning techniques, particularly Monte-Carlo Tree Search (MCTS) and reinforcement learning, have emerged as useful additions, significantly expanding the capabilities of modern planning systems.
With the advent of powerful large language models (LLMs), there are new opportunities to both revisit classical planning problems, and to further explore new problems through natural language specification that reflects the ambiguity and uncertainty of real-world domains. Planning capability is important for many tasks such as game playing, meeting scheduling and trip planning. Research is already underway to leverage the commonsense knowledge of LLMs in real-world tasks (Huang et al., 2022; Singh et al., 2023; Ding et al., 2023) and to generate plans (Valmeekam et al., 2023; Hao et al., 2023; Guan et al., 2024). The research has shed some light on LLMs’ struggle with planning tasks. Even state-of-the-art LLMs may produce ineffective or even incorrect plans, even in straightforward scenarios. Our paper focuses on analyzing and improving the planning capability of LLM systems.
We provide a scalable benchmark suite in both PDDL and natural language to measure planning capability of LLMs. Specifically, we explore two distinct planning representations: the formal Planning Domain Definition Language (PDDL) (McDermott et al., 1998), which provides a standardized representation for classical planning problems and allows for rigorous plan validation; and natural language, which offers a more flexible and intuitive representation better reflecting real-world scenarios. For both scenarios, we provide a code for generating as many instances with a degree of difficulty of choice. We also provide a mapping method for translating PDDL benchmarks to natural language and measure the performance of the generated benchmarks. The generated planning tasks are scalable and can grow to examine and assist stronger models. Figure 1 depicts a simple sample of BlocksWorld benchmark with description both in PDDL and natural language. Each instance in BlocksWorld consists of a set of blocks, a table and a robot hand, where the goal is to move from one block configuration to another. The problem definition includes list of objects, initial and goal states. The figure also includes the LLM’s output plan from the in-context learning scenario, which we extensively analyze using various LLM models and benchmarks in Figure 2.
# PDDL
(define (problem BW)
(:domain bw-4ops)
(:objects A B C) (:init
@@
# Planning: LLM generates PDDL.
@@
The goal is:
C is on B.
A is on C.
# Planning: LLM generates NL plan
@@
Figure 1: A simple instance of BlocksWorld planning benchmark: visual, PDDL and natural language descriptions side by side. Problem definition includes list of objects, initial and goal states. We have also included LLM’s output plan in the in-context learning scenario of Figure 2.
@@
We explore the planning capability of LLMs using both In Context Learning (ICL) through the many-shot paradigm as well and through chain-of-thought (CoT) (Wei et al., 2022) inference time techniques; we also explore fine-tuning strategies. We observe that carefully instructing the model using ICL leads to a significant boost in planning performance, which can be further improved by using a many-shot approach with long context. Moreover, CoT reasoning strategies (MCTS, Tree-of-thought, debate-as-reasoning) allow smaller models to perform closer to SoTA frontier models in the natural language task domain. Our results show that fine-tuning with the optimal plan can lead to near-perfect accuracy, even when using relatively small models.
Next, we investigate both in-domain and out-of-domain generalization and note that the proposed plans demonstrate excellent in-domain generalization: instances with similar complexity are solved with similar levels of accuracy. For ICL, prompting with easier instances leads to better performance on hard instances compared to prompting with hard instances. Fine-tuning leads to better performance for both in-domain and out-of-domain scenarios, even with a much smaller model, compared to ICL with long-context.
Lastly, we probe the failure modes of the model. We categorize the failure modes into three categories: failure to satisfy environmental constraints, failure to meet the goal and failure to generate legal actions in a given state. We note that not all of these failure modes are present across all benchmarks and methods. Moreover, as the instance complexity increases the model success rate decreases. Additionally, we pinpoint failure modes that are result of biases in training data emphasizing the importance of data curation during training.
@@ 1.1 Related Work
Prior works investigations the planning capabilities of LLMs, have found that these models struggle to solve planning tasks (Hao et al., 2023; Valmeekam et al., 2023; 2024; Kambhampati et al., 2024), In contrast, we show that LLMs can be capable of solving such tasks and one can reach near-perfect accuracy for some scenarios, with certain methods. We run experiments on the same BlocksWorld benchmark as these prior works, and generate additional difficult cases (i.e., adding more blocks). Similar to these works, we also use PDDL verifiers to compute accuracy. See Section 3.4 for details.
Xie et al. (2024) proposed a TravelPlanner benchmark and showed GPT4-Turbo can solve some of the benchmark tasks with a success rate of $0.6\%$. The TripPlanner benchmark used in our work has two main differences: it is not an agent-based environment and rather a natural language benchmark, and it has unique answers due to carefully designed constraints.
Stechly et al. (2024) suggests that LLMs are not capable of generalizing their plans on BlocksWorld if one uses chain-of-thought (CoT) (Wei et al., 2022). In this work we show positive results in terms of generalization performance.
@@
In addition to standard few-shot prompting, we also make use of inference time techniques that construct a chain-of-thought (Wei et al., 2022), namely tree-of-thought (ToT) (Yao et al., 2023) and Monte-Carlo Tree Search (MCTS) (Hao et al., 2023) and debate-as-reasoning (Du et al., 2023). We demonstrate that these methods can considerably improve an LLMs planning capabilities such that smaller open-source models that use these approaches can outperform larger foundational models.
2 Benchmark
To evaluate the planning capabilities of LLMs, we assemble a benchmark suite that appropriately represents various classical and natural-language planning tasks. On these benchmarks, we assess LLM performance using In-Context Learning (ICL), Supervised Fine-Tuning (SFT), and chain-of-thought (CoT) methods for planning.
PDDL (Planning Domain Definition Language): PDDL McDermott et al. (1998) is a standardized language used in artificial intelligence for representing planning problems. PDDL provides a formal way to describe the initial state of an environment, a goals, the space of valid actions, and the state-transition properties of actions in the environment. PDDL has two main components (1) Domain: Describes the general characteristics of the planning problem, including the types of objects, actions, and predicates (conditions that can be true or false). (2) Problem: Defines a specific instance of the planning problem within the domain, including the initial state of the world and the goals to be achieved.
We select three datasets that use PDDL, generated additional subsets of them, and additionally map these datasets to natural language for an additional evaluation task. Additionally, we select two native natural language datasets, containing Trip Planning and Calendar Scheduling tasks (Zheng et al., 2024). For the PDDL-based datasets, we select BlocksWorld, Logistics, and Mini-Grid. We then translated the PDDL problem descriptions from these datasets into natural language to compare performance when using formal and informal representations.
2.1 PDDL Benchmarks
@@
This procedure is repeated with increasingly difficult configurations for a select number of problems. The result of this procedure are additional datasets that comprise a set of PDDL problems and solutions of various difficulty. Importantly, this procedure enables us to create datasets with increasingly difficult problems and any number of samples, which are appropriate for assessing the ability to plan using different methods, such as in-context learning versus Supervised Fine-Tuning. Moreover, we can scale the dataset generation and create as many instances as needed for different investigations. We provide the details of our benchmark suite below.
@@
BlocksWorld: BlocksWorld is a standard planning problem from International Planning Conference (IPC)-2000  [^2] https://github.com/potassco/pddl-instances/tree/master/ipc-2000. This domain consists of a set of blocks, a table and a robot hand, where the goal is to move from one block configuration to another. We generate a dataset for 3 to 7 blocks. As detailed in the Appendix B.1, we produced 28k unique BlocksWorld samples. From these, 25.5k were randomly selected for the training set and 2,500 for the validation set.
Logistics: Logistics is an AI planning problem from IPC-1998  [^3] https://github.com/potassco/pddl-instances/tree/master/ipc-1998 expressed in PDDL that involves arranging the delivery of packages to their destinations using trucks within cities and airplanes between cities. The aim is to optimize transportation modes under constraints such as vehicle capacities and locations, showcasing model’s ability to manage multi-step logistics efficiently.
Mini-Grid: Mini-Grid is a task from Artificial Intelligence Planning Systems (AIPS)-1998  [^4] https://github.com/AI-Planning/pddl-generators/tree/main/minigrid, also expressed in PDDL. We create various floorplans with rooms containing random configurations of key shapes. The goal then is for a robot to navigate from an initial position to a designated goal cell.
@@ 2.2 Native Natural Language Planning Benchmarks
Trip Planning: Trip Planning is a task from NaturalPlan (Zheng et al., 2024) benchmark focusing on planning a trip itinerary under given constraints. The goal of the task is to find an itinerary satisfying constraints such as the order of visiting N cities. It includes enough constraints for each instance such that there is only one solution to the task, which makes the evaluation of the predictions straightforward.
Calendar Scheduling: Calendar Scheduling from the NaturalPlan (Zheng et al., 2024) benchmark represents the task of scheduling a meeting of either 30 minutes or an hour among up to 7 attendees. The attendees may have a busy schedule or a light schedule with less than half of the working hours spent in meetings.
@@ 3 Experiments
Previous works demonstrated that, without intervention, LLMs often struggle with even simple planning tasks (Hao et al., 2023; Valmeekam et al., 2023; 2024; Kambhampati et al., 2024). LLMs often lack the information on how to structure their plan constructively, and struggle to plan around the enumerated constraints. In this section we present our experimental results for the interventions we investigate, demonstrating that they lead to significant improvements to the LLMs planning capability. We also investigate plan generalization (i.e., the ability to generalize to unseen instances) in several scenarios.
For PDDL experiments, we measure accuracy of the generated plan with a verifier (Howey et al., 2004). For natural language experiments we rely on either recasting the task to PDDL and verifying with a verifier, or extracting the answer and comparing it to expected results (Zheng et al., 2024). In all experiments, GPT-4 refers to GPT-4 Turbo and we may omit "Turbo" for space constraints, also if we do not mention Pro or Flash Gemini 1.5 refers to Gemini 1.5 Pro.
@@
Figure 2: PDDL Planning and Natural Language Planning with few-shots for different LLMs. Natural language text are generated from formal PDDL problem definitions.
3.1 In-context learning
For in-context learning (Brown et al., 2020), we adhere to the standard procedure, employing a prompt containing several examples for the task. Each example comprises a planning problem statement and its corresponding solution, referred to as a shot. Following the examples, the test problem is added without the corresponding solution. Subsequently, an LLM receives this prompt and is expected to generate a plan following the format and logic of the examples in the prompt. See Appendix A for examples of prompts.
3.1.1 Many-shot in-context learning
We evaluate the planning capability of the model as we add more examples (“shots") into the context, inspired by the success of many-shot learning across a large number of tasks (Agarwal et al., 2024). The challenge of “in-context planning" involves understanding a specific task and problem through a limited number of examples. Additionally, it requires the models to produce a solution without checking each planning step to confirm if a proposed move is correct. The model has to create a plan in a single inference step, keeping ‘in mind’ all the constraints the task imposes.
Figure 2 shows the in-context learning performance on classical planning and natural language benchmarks as we vary the number of shots. We consider Gemini 1.5 Pro (GeminiTeam et al., 2024b), GPT-4 Turbo (OpenAI et al., 2024), Gemini 1.5 Flash (GeminiTeam et al., 2024b) and Gemma 2 27b (Team et al., 2024) models. Overall, we notice that for natural language and PDDL scenarios, models have similar trends in terms of planning accuracy as we increase the number of shots. Moreover, different models are impacted differently as we provide additional number of shots, e.g., Gemini 1.5 Pro outperforms other models both in one shot scenario and as we increase the number of shots; indicating that the model not only can plan better with a fewer number of examples/shots, it can also make effective use of additional and longer context. Gemini 1.5 Flash -a smaller, faster and more efficient model than Gemini 1.5 Pro is generally outperformed by Gemini 1.5 Pro but occasionally matches GPT-4 performance.
BlocksWorld: Figure 2(a)2(d) show the performance of Gemini 1.5 models on this benchmark as we increase the number of few-shot examples. We note that as we increase the number of shots GPT-4 performance increases while Gemini 1.5 Pro’s performance saturates or degrades as we go beyond 40 shots. The 1-shot planning capability of Gemini 1.5 Pro and Gemini 1.5 Flash reaches reaches 35% and 26%, while GPT-4 performance is close to zero. Moreover the 40-shots planning capability of Gemini 1.5 Pro reaches 48% range which performs better than the best (200-shots) performance of GPT-4, which peaks at 43%.